OpenStudy (anonymous):

log4(x+16)=3 4 is at the foot of log

OpenStudy (anonymous):

\[\log_{4}(x+16)=3 \]

OpenStudy (anonymous):

x+16=64 x=48

OpenStudy (anonymous):

thanks but could you show how you got there I am horrible at this

OpenStudy (anonymous):

http://www.twiddla.com/583183

OpenStudy (lalaly):

x+16 = 4^3 x+16 = 64 x=48

OpenStudy (anonymous):

come to twiddla